OpenStudy (anonymous):

Please Help! Find the real solutions of the equation by graphing. x^2-x+2=0

OpenStudy (phi):

do you have a graphing calculator?

OpenStudy (anonymous):

No i dont

OpenStudy (phi):

I would make up a table of x y pairs pick some numbers like x= -3 up to +3, and find the y once you get a list of (x,y) points, plot them. Can you find the list of (x,y) points (as a first step) ??

OpenStudy (anonymous):

well i tryed to solve the question like this x^2-x+2=0 (x-2)(x-1)=0 x=2 x=1 could I use these numbers?

OpenStudy (anonymous):

in this question , no solution.

OpenStudy (anonymous):

@ASAAD123 What do you mean it has it be a graph? What do you mean no solution

OpenStudy (phi):

no, because (x-2)(x-1) multiplies out to x^2 -3x +2 It is not your equation. try finding some (x,y) pairs.

OpenStudy (anonymous):

No real roots. Curve does not cross the x axis. Refer to the attachment.

OpenStudy (anonymous):

@robtobey So its x=3

OpenStudy (anonymous):

\[x=\frac{1}{2} \left(1\pm i \sqrt{7}\right) \]

OpenStudy (anonymous):

I think I got it Thanks everyone! I'm not sure who to give a medal too?

OpenStudy (anonymous):

the curve does not touch and pass through x-axis.

OpenStudy (anonymous):

I don't want to ruffle any one's feathers on this site, however, in some circles this problem would be viewed as a "trick question".

OpenStudy (anonymous):

Could someone help with this one Divide using synthetic division -3x^3+11x^2-x+20by(x-4)

OpenStudy (anonymous):

\[-3 x^2-x-5 \]with no remainder.

OpenStudy (anonymous):

This one says to use synthetic division and the given factor to completly factor the polynomial

OpenStudy (anonymous):

Well the remainder is 120. Your job is to verify that the quotient is\[x^2+5 x+30 \]using synthetic division.

OpenStudy (anonymous):

in the question 24x?

OpenStudy (anonymous):

@ASAAD123 whats do you mean? Whats your question?

OpenStudy (anonymous):

in the question,you write 24 without x.

OpenStudy (anonymous):

you are right its x^3-x^2-24x-36;(x-6)

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Thanks soooo much!

OpenStudy (anonymous):

welcome